Companies invest in expansion projects with the expectation of increasing the eamings of its business. Consider the case of Ghelp me!
Answer #1

let me know if you need any clarification..

Ans a Computation of NPV using Acceperated depreciation method
Year -0 Year -1 Year -2 Year -3 Year -4
initial investment -15000
i unit sales           3,500           4,000           4,200           4,250
ii Sales price          38.50          39.88          40.15          41.55
iii variable cost          22.34          22.85          23.67          23.87
iv=ii-iii Contribution per unit          16.16          17.03          16.48          17.68
v=i*iv Total contribution        56,560        68,120        69,216        75,140
vi Fixed cost        37,000        37,500        38,120        39,560
vii Depreciation rate 33% 45% 15% 7%
viii Depreciation amount           4,950           6,750           2,250           1,050
ix=v-vi-viii Profit before tax        14,610        23,870        28,846        34,530
x=ix*40% Tax @ 40%           5,844           9,548        11,538        13,812
xi=ix-x Profit after tax           8,766        14,322        17,308        20,718
xii=xi+viii Operating cash flow        13,716        21,072        19,558        21,768
xiii=xii+viii Total cash flow -15000        13,716        21,072        19,558        21,768
xiv PVIF @ 11%     1.0000        0.9009        0.8116        0.7312        0.6587
xv=xiii*xiv present value -15000        12,357        17,103        14,300        14,339        43,099
Therefore NPV = 43099
Ans b Computation of NPV using stratight line depreciation
Year -0 Year -1 Year -2 Year -3 Year -4
initial investment -15000
Total contribution              -          56,560        68,120        69,216        75,140
Fixed cost        37,000        37,500        38,120        39,560
Depreciation amount           3,750           3,750           3,750           3,750
Profit before tax        15,810        26,870        27,346        31,830
Tax @ 40%           6,324        10,748        10,938        12,732
Profit after tax           9,486        16,122        16,408        19,098
Operating cash flow        13,236        19,872        20,158        22,848
Total cash flow    (15,000)        13,236        19,872        20,158        22,848
PVIF @ 11%     1.0000        0.9009        0.8116        0.7312        0.6587
present value    (15,000)        11,924        16,129        14,739        15,051        42,843
Ans c using Accelerated depreciation method will result in the hiehgest NPV for the project
Ans d Year -1 Year -2 Year -3 Year -4
i reduction in after tax cash flow 500 500 500 500
ii PVIF @ 11%     0.9009        0.8116        0.7312        0.6587
iii=i*ii present value           450              406              366              329           1,551
Correct answer =        1,551
Ans e Correct answer is option : Increase the amount of the initial investment by $9000